Bear population. To estimate the size of the bear population on the Keweenaw Peninsula, conservationists captured, tagged, and released 50 bears. One year later, a random sample of 100 bears included only 2 tagged bears. What is the conservationist’s estimate of the size of the bear population?
Answer by stanbon(75887)

Bear population. To estimate the size of the bear population on the Keweenaw Peninsula, conservationists captured, tagged, and released 50 bears. One year later, a random sample of 100 bears included only 2 tagged bears. What is the conservationist’s estimate of the size of the bear population?
-----------------------------------------
(population #)/(# caught and tagged) = (# caught including tagged)/(# tagged)
--------
(population #)/50 = 100/2
population # = 50(100/2)
population estimate = 2500 bears
============================
Cheers,
Stan H.